The length of a rectangle is 4 less than 3 times the width. If the perimeter is 40, find the dimensions of the rectangle. What is this answer??
Answer by awang1996(50)

The length of a rectangle is 4 less than 3 times the width. If the perimeter is 40, find the dimensions of the rectangle. What is this answer??


3w-4=l
3w-4+3w-4+w+w=40
8w-8=40
8w=48


w=6
l=14